A man enters a tall tower, needing to know its height. He notes that a long pendulum extends from the ceiling almost to the floor and that its period is 14.0 s.

(a) How tall is the tower?
m

(b) If this pendulum is taken to the Moon, where the free-fall acceleration is 1.67 m/s2, what is the period there?
s
Answer #1

Part A)

Apply the following

T = 2\pi \sqrt{\frac{l}{g}}

14 = 2(3.14)sqrt(L/9.8)

L = 48.7 m

Part B)
Same formula, different g

T = 2(3.14)sqrt (48.7)/1.67)

T = 33.9 sec
Answer #2

A)
tall of the tower is L = T^2*g/(4*pi^2) = 14^2*9.8/(4*3.142^2) = 48.65m...
B)T = 2*pi*sqrt(L/g) = 2*3.142*sqrt(48.65/1.67) = 33.92 sec